OpenStudy (anonymous):

Given the literal equation ab/a+b=c with b = –1 + a, what is the value of c if a = 5? a. 20/9 b. 25/9 c. 4 d. 11

OpenStudy (anonymous):

a

OpenStudy (texaschic101):

.b = -1 + a a = 5 b = -1 + 5 b = 4 ab / (a + b) = c (5)(4) / (5 + 4) = c 20/9 = c answer is A....

OpenStudy (anonymous):

Thank You !